# A Bayesian latent Gaussian conditional autoregressive copula model for analyzing spatially-varying trends in rainfall

### v1

@@ 2.3 Copulas
Copulas provide a flexible approach to modeling and estimating the joint distribution of variables independently of their marginal distributions, making them useful for capturing complex dependencies. A copula is a joint probability distribution function that ensures both marginal distributions are $\textrm{Uniform}(0,1)$. Specifically, a function $C(y_{1},y_{2})$ qualifies as a copula if $C(0,0)=0$ and for $y_{1},y_{2}\in(0,1)$, it satisfies $C(y_{1},1)=y_{1},\ C(1,y_{2})=y_{2}.$ Suppose we aim to determine a suitable joint probability distribution function $H(y_{1},y_{2})$ for random variables $Y_{1}$ and $Y_{2}$, given that their marginal distributions follow the continuous distribution functions $F_{1}$ and $F_{2}$, respectively. That is, given $F_{1}(y_{1})$ and $F_{2}(y_{2})$, along with knowledge of the dependency structure between $Y_{1}$ and $Y_{2}$, we seek to define an appropriate joint distribution function as $H(y_{1},y_{2})=P(Y_{1}\leq y_{1},Y_{2}\leq y_{2})$. Since the transformed variables $U_{1}=F_{1}(Y_{1})$ and $U_{2}=F_{2}(Y_{2})$ follow a uniform distribution on $(0,1)$, the joint distribution function of $U_{1}$ and $U_{2}$ forms a copula. Then
@@
In our context, we consider a random vector following a standard bivariate Gaussian distribution with correlation $\rho$. The Gaussian copula function is defined as $C_{\rho}(u,v)=\Phi_{\rho}(\Phi^{-1}(u),\Phi^{-1}(v))$, where $u,v\in[0,1],\Phi(\cdot)$ represents the standard normal cumulative distribution function (CDF), and $\Phi_{\rho}(\cdot,\cdot)$ denotes the CDF of a standard bivariate Gaussian distribution with correlation $\rho$. To construct a bivariate Gamma distribution, we assume that the random vector $(Y_{1},Y_{2})$ has gamma-distributed marginals given by the distribution functions $F_{1}$ and $F_{2}$, respectively. Using the Gaussian copula, we define $(Y_{1},Y_{2})=\left(F_{1}^{-1}(\Phi(X_{1})),F_{2}^{-1}(\Phi(X_{2}))\right)$, where $F_{1}^{-1}$ and $F_{1}^{-1}$ are the marginal quantile functions for $Y_{1}$ and $Y_{2}$, and $(X_{1},X_{2})\sim\Phi_{\rho}$. The joint CDF of $Y_{1}$ and $Y_{2}$ is then given by $H(Y_{1},Y_{2};\rho)=C_{\rho}(\Phi(X_{1}),\Phi(X_{2})),$ while ensuring that the marginal distributions of $Y_{1}$ and $Y_{2}$ remain $F_{1}$ and $F_{2}$, respectively.
An appropriate copula should accurately represent the assumed dependencies between $U_{1}$ and $U_{2}$. Since $F_{1}$ and $F_{2}$ are increasing functions, the dependence structure induced by the chosen copula should reflect the relationship we expect between $Y_{1}$ and $Y_{2}$. For example, if we assume that $Y_{1}$ and $Y_{2}$ have a correlation of $\rho$, we should select a copula such that the random variables following this copula exhibit a correlation of $\rho$. However, since correlation captures only linear dependencies, the correlation between $Y_{1}$ and $Y_{2}$ does not necessarily match the correlation between $U_{1}$ and $U_{2}$ [51].
@@ 2.4 Uniform-transformed Quantile-Quantile (QQ) plot
Model validation is essential to ensure the reliability of statistical inference. Suppose we are interested in fitting a marginal distribution model $F(\cdot,\bm{\theta}_{i})$ where $\bm{\theta}_{i}$ denotes the model parameter corresponding to the $i$th observation $Y_{i}$ and possibly different across observations that we can simplify in terms of some covariates under a regression model setup. Based on some estimates of $\bm{\theta}_{i}$, say $\widehat{\bm{\theta}}_{i}$, approximately $Y_{i}\sim F(\cdot,\widehat{\bm{\theta}}_{i})$ and thus $F(Y_{i},\widehat{\bm{\theta}}_{i})\sim\textrm{Uniform}(0,1)$ if the model $F$ is appropriate. We can thus consider a QQ plot of low through high quantiles of the $\textrm{Uniform}(0,1)$ distribution and the corresponding quantiles of the empirical CDF of $F(Y_{i},\widehat{\bm{\theta}}_{i})$s as a diagnostic tool. We call this tool a Uniform-transformed QQ plot [52]. If the plot remains close to the $y=x$ line, we can safely consider $F$ a reasonable model. If we compare two models, $F_{1}$ and $F_{2}$, we can calculate and compare the uniformed-transformed QQ plots for both distributions. While we can use a Kolmogorov-Smirnov test for a similar purpose, the uniformed-transformed QQ plots showcase the range of quantiles where the model fitting is reasonable; for example, a model may fit the data distribution well in the bulk and poorly in the tails, and this tool can explain the scenario.
@@ 2.5 Moran’s I for assessing spatial autocorrelation
Areal data often exhibit autocorrelation, meaning that observations at nearby, possibly neighboring, regions are more similar than distant ones. Moran’s I [53] is a widely used measure for detecting spatial autocorrelation, and it is calculated as
@@
where $Y_{i}$ represents the observed value at region $i$, $w_{ij}$ is the spatial weight matrix defining neighborhood relationships in (4), $\bar{Y}$ is the mean value of $Y_{i}$, and $N$ is the total number of locations. Moran’s I ranges from $-1$ to $+1$, where positive values of Moran’s I suggest clustering, negative values indicate dispersion, and values near zero suggest spatial randomness. Under the null hypothesis of complete randomness, $\text{E}(I)=-1/(N-1)$ and $\text{Var}(I)$ can be calculated in terms of $w_{ij}$s, and we thus can obtain a $z$-score by centering and scaling. Finally, we obtain the $p$-value by assuming normality for the test statistic and decide on the spatial randomness accordingly.
@@ 2.6 Bayesian model comparison
To compare multiple Bayesian models with the same likelihood structure but different prior structures, we report Deviance information criteria [DIC, 54] and Watanabe-Akaike information criteria [WAIC, 55]. We calculate DIC as
$\textrm{DIC}=\bar{D}+p_{D}=D(\bm{Y}|\widehat{\bm{\theta}})+2p_{D},$
where $D(\bm{Y}|\bm{\theta})=-2\log\left[p(\bm{Y}|\bm{\theta})\right]$ is the deviance for measuring model fit, $\bar{D}$ is the posterior mean deviance defined as $\text{E}(D(\bm{Y}|\bm{\theta})|\bm{Y})$, $\widehat{\bm{\theta}}=\text{E}(\bm{\theta}|\bm{Y})$ is the posterior mean of the model parameter vector $\bm{\theta}$, and $p_{D}=\bar{D}-D(\bm{Y}|\widehat{\bm{\theta}})$ is the effective number of parameters in the model. Given the MCMC outputs, calculating DIC is straightforward. The intuition is that models with smaller DIC are parsimonious (small $p_{D}$) and fit well (small $\bar{D}$).
Further, we can compute WAIC, also known as the widely applicable information criteria, as an alternative to DIC. Here, WAIC approximates $n$-fold (i.e., leave-one-out) cross-validation. Suppose we have $n$ independent observations (possibly vectors) $\bm{Y}_{i}$ for $i=1,\ldots,n$ and $\bm{\theta}$ is the parameter vector in the model (likelihood) $f(\cdot)$. Let $m_{i}$ and $v_{i}$ be the posterior mean and variance of $\log\left[f\left(\bm{Y}_{\bm{i}}|\bm{\theta}\right)\right]$. The effective model size is $p_{W}=\sum_{i=1}^{n}v_{i}$ and WAIC is given by
@@
The measure DIC is computationally simpler than WAIC but may be unreliable for hierarchical and complex models. At the same time, WAIC is more general and provides a better approximation of predictive performance, making it preferable in most Bayesian applications. [56] recommend using WAIC or leave-one-out cross-validation over DIC for model comparison in modern Bayesian workflows.
@@ 3 Data description and exploratory analysis
Unlike the administrative state-wise boundaries, mainland India (excluding Andaman and Nicobar Islands and Lakshadweep Archipelago) comprises 34 meteorological subdivisions; the division is determined based on meteorological homogeneity [4]. We present the geographical boundaries of these 34 subdivisions in Figure 2. We obtain annual total rainfall data (in mm) for these subdivisions, covering the years 1951 to 2014, from the Open Government Data (OGD) Platform, India (https://data.gov.in), along with shape files of the boundaries. Based on the rainfall data collected from 641 districts across India by the India Meteorological Department (IMD), [4] computed monthly rainfall amounts for these districts by averaging the rainfall data from available stations within each district for each month. An area-weighted averaging method was then applied to derive subdivision-wise rainfall data from the district-wise data. [57] provide a more detailed report on the dataset preparation. The rest of the paper focuses on analyzing this dataset, which we refer to as the “Indian rainfall dataset”.
Suppose we denote the annual total rainfall for the $i$th meteorological subdivision and $t$th year by $Y_{it}$, for $i=1,\ldots,n$ and $t=1,\ldots,T$ with $n=34$ and $T=64$. Given the extensive $64$–year observational period encompassing phases of global warming, significant trends across the years are plausible. The data are positive-valued; hence, we can model the marginal distributions using log-normal or gamma regression, for example, where we can model the mean as a function of the year. A log-normal distribution assumption is more popular in general spatial or areal data analysis literature because we can consider a logarithmic transformation of the data and model them using an ordinary or spatial linear regression framework instead of a generalized ordinary or spatial linear regression framework. At the same time, as mentioned in Section 1, the gamma distribution assumption for the margins is the most popular in rainfall modeling literature.
@@
We first postulate that the data are independently distributed across years and marginally follow a gamma distribution. To model the space-time varying mean parameter, we write $\text{E}(Y_{it})=\mu_{it}$ with areally-varying shape parameter $a_{i}$. Hence, for every region index $i$, we assume
@@
where $\alpha_{i}$ is the shape parameter and $\mu_{it}$ is the mean of the gamma distribution for region index $i$ and time point $t$, which is connected with a log-link function
@@
in which $\alpha_{i}$ and $\beta_{i}$ are areally-varying coefficients. A competing log-normal model would be $Y^{*}_{it}=\log(Y_{it})\overset{\mathrm{ind}}{\sim}\text{Normal}(\mu_{it},\sigma^{2}_{i}),\ t=1,\dots,T$, where $\mu_{it}=\alpha_{i}^{*}+\beta_{i}^{*}t$ for some areally-varying coefficients $\alpha^{*}_{i}$, $\beta^{*}_{i}$, and variances $\sigma^{2}_{i}$. We separately fit the regression models for every region index $i$ using maximum likelihood estimates and compare their uniform-transformed QQ plots in the left panel of Figure 1. The lognormal model better fits low quantile levels, while the gamma model outperforms the lognormal model overall. Compared to the lognormal model, the gamma model provides approximately 7% lower root mean square error and mean absolute error (in terms of fitting). For every subregion $i$, we present the uniform-transformed quantiles based on gamma regression in the right panel of Figure 1. Except for certain middle-range quantiles, the model fits the data reasonably well. In the rest of the paper, we subsequently consider the gamma model in (7) and (8).
@@
We further reparameterize the model in (7) and (8) in terms of a standard representation of a gamma distribution using shape and rate parameters as follows. First we rewrite (8) as $\log(\mu_{it})=\tilde{\alpha}_{i}+\tilde{\beta}_{i}t^{*}$, where $\tilde{\alpha}_{i}=\alpha_{i}+\beta_{i}m_{t}$, $\tilde{\beta}_{i}=\beta_{i}s_{t}$, and $t^{*}=(t-m_{t})/s_{t}$ with $m_{t}=T^{-1}\sum_{t=1}^{T}t$ and $s_{t}^{2}=T^{-1}\sum_{t=1}^{T}t^{2}-m_{t}^{2}$. Then, following (7), the rate parameter is given by
$\lambda_{it}=\frac{a_{i}}{\mu_{it}}=a_{i}\exp[-\tilde{\alpha}_{i}]\exp[-\tilde{\beta}_{i}t^{*}]=a_{i}b_{i}\exp[c_{i}t^{*}],$ (9)
where $c_{i}$, the coefficient of $t^{*}$, captures the climate change behavior over the years for the meteorological subdivision $i$. In this parameterization, the mean and variance of $Y_{it}$ can be obtained as
$\begin{split}\text{E}(Y_{it})&=\mu_{it}=\frac{a_{i}}{\lambda_{it}}=\frac{a_{i}}{a_{i}b_{i}\exp[c_{i}t^{*}]}=b_{i}^{-1}\exp[-c_{i}t^{*}]=\exp[\alpha_{i}+\beta_{i}t],\\ \text{Var}(Y_{it})&=\frac{a_{i}}{\lambda_{it}^{2}}=\frac{a_{i}}{a_{i}^{2}b_{i}^{2}\exp[2c_{i}t^{*}]}=\frac{1}{a_{i}}(b_{i}^{-1}\exp[-c_{i}t^{*}])^{2}=\frac{1}{a_{i}}\mu_{it}^{2}.\end{split}$ (10)
Here, the coefficient of variation, $a_{i}^{-1/2}$ for region $i$, remains constant over $t$. We further interpret $a_{i}$ as a shape parameter, $b_{i}$ as an intercept-related parameter, and $c_{i}$ as a slope-related parameter; the rest of the paper follows this parameterization.
For exploratory analysis, we fit the model in (7) and (9) separately for every region $i$ using maximum likelihood estimation procedure and we present the estimates (MLEs) of $\bm{a}=(a_{1},\dots,a_{n})^{\top}$, $\bm{b}=(b_{1},\dots b_{n})^{\top}$ and $\bm{c}=(c_{1},\dots,c_{n})^{\top}$ in Figure 2. The standard errors of the estimates are low due to $T=64$ replications used for the analysis, and hence, any spatial pattern visible in Figure 2 can be conceptualized to build a flexible model for them. First, we notice that the MLEs of $a_{i}$s, $b_{i}$s, and $c_{i}$s are non-constant across space, motivating us to consider areally-varying coefficients. Second, the estimates for nearby regions exhibit similarity, indicating the suitability of a CAR or an ICAR prior for the parameters.

### v2

@@ 2.3 Copulas
Copulas offer a flexible approach to modeling and estimating the joint distribution of variables, independent of their marginal distributions, making them useful for capturing complex dependencies [sklar1959distribution, nelsen2006introduction]. A copula is a joint probability distribution function that ensures both marginal distributions are $\textrm{Uniform}(0,1)$. Specifically, a function $C(y_{1},y_{2})$ qualifies as a copula if $C(0,0)=0$ and for $y_{1},y_{2}\in(0,1)$, it satisfies $C(y_{1},1)=y_{1},\ C(1,y_{2})=y_{2}.$ Suppose we aim to determine a suitable joint probability distribution function $H(y_{1},y_{2})$ for random variables $Y_{1}$ and $Y_{2}$, given that their marginal distributions follow the continuous distribution functions $F_{1}$ and $F_{2}$, respectively. That is, given $F_{1}(y_{1})$ and $F_{2}(y_{2})$, along with knowledge of the dependency structure between $Y_{1}$ and $Y_{2}$, we seek to define an appropriate joint distribution function as $H(y_{1},y_{2})=P(Y_{1}\leq y_{1},Y_{2}\leq y_{2})$. Since the transformed variables $U_{1}=F_{1}(Y_{1})$ and $U_{2}=F_{2}(Y_{2})$ follow a uniform distribution on $(0,1)$, the joint distribution function of $U_{1}$ and $U_{2}$ forms a copula. Then
@@
In our context, we consider a random vector following a standard bivariate Gaussian distribution with correlation $\rho$. The Gaussian copula function is defined as $C_{\rho}(u,v)=\Phi_{\rho}(\Phi^{-1}(u),\Phi^{-1}(v))$, where $u,v\in[0,1],\Phi(\cdot)$ represents the standard normal cumulative distribution function (CDF), and $\Phi_{\rho}(\cdot,\cdot)$ denotes the CDF of a standard bivariate Gaussian distribution with correlation $\rho$. To construct a bivariate Gamma distribution, we assume that the random vector $(Y_{1},Y_{2})$ has gamma marginal distributions given by the distribution functions $F_{1}$ and $F_{2}$, respectively. Using the Gaussian copula, we define $(Y_{1},Y_{2})=\left(F_{1}^{-1}(\Phi(X_{1})),F_{2}^{-1}(\Phi(X_{2}))\right)$, where $F_{1}^{-1}$ and $F_{1}^{-1}$ are the marginal quantile functions for $Y_{1}$ and $Y_{2}$, and $(X_{1},X_{2})\sim\Phi_{\rho}$. The joint CDF of $Y_{1}$ and $Y_{2}$ is then given by $H(Y_{1},Y_{2};\rho)=C_{\rho}(\Phi(X_{1}),\Phi(X_{2})),$ while ensuring that the marginal distributions of $Y_{1}$ and $Y_{2}$ remain $F_{1}$ and $F_{2}$, respectively.
An appropriate copula should accurately represent the assumed dependencies between $U_{1}$ and $U_{2}$. Since $F_{1}$ and $F_{2}$ are increasing functions, the dependence structure induced by the chosen copula should reflect the relationship we expect between $Y_{1}$ and $Y_{2}$. For example, if we assume that $Y_{1}$ and $Y_{2}$ have a correlation of $\rho$, we should select a copula such that the random variables following this copula exhibit a correlation of $\rho$. However, since correlation captures only linear dependencies, the correlation between $Y_{1}$ and $Y_{2}$ does not necessarily match the correlation between $U_{1}$ and $U_{2}$ [ross2022simulation].
While the above exposition uses two random variables for clarity, copulas extend naturally to higher dimensions. For a random vector $(Y_{1},Y_{2},\dots,Y_{n})^{\intercal}$ with continuous marginal distributions $F_{1},F_{2},\dots,F_{n}$, Sklar’s theorem [sklar1959distribution] suggests the existence of a unique copula $C$ such that for joint CDF $H$ of $(Y_{1},Y_{2},\dots,Y_{n})$, we have
$\begin{split}H(y_{1},y_{2},\dots,y_{n})&=P(Y_{1}\leq y_{1},Y_{2}\leq y_{2},\dots,Y_{n}\leq y_{n})\\ &=P(F_{1}(Y_{1})\leq F_{1}(y_{1}),F_{2}(Y_{2})\leq F_{2}(y_{2}),\dots,F_{n}(Y_{n})\leq F_{n}(y_{n}))\\ &=C(F_{1}(y_{1}),F_{2}(y_{2}),\dots,F_{n}(y_{n})).\end{split}$ (6)
While the above generalization for any dimension $n$ is possible, only a few classes of copulas, like the Gaussian or the Student-$t$ copula, can be easily computed when $n$ is large. In our application, we employ a Gaussian copula, where we parameterize the dependence through the correlation (or precision) matrix that reflects the neighborhood structure of the Indian meteorological subdivisions.
@@ 2.4 Uniform-transformed Quantile-Quantile (QQ) plot
Model validation is essential to ensure the reliability of statistical inference. Suppose we aim to fit a marginal distribution $F(\cdot,\bm{\theta}_{i})$ for each observation $Y_{i}$, where the parameter $\bm{\theta}_{i}$ may vary across $i$. One can often explain such variation through covariates under a regression framework. Based on estimates of $\bm{\theta}_{i}$, say $\widehat{\bm{\theta}}_{i}$, approximately $Y_{i}\sim F(\cdot,\widehat{\bm{\theta}}_{i})$ and thus by the probability integral transform, $F(Y_{i},\widehat{\bm{\theta}}_{i})\sim\textrm{Uniform}(0,1)$ if the model $F$ is appropriate, meaning that it correctly represents the true marginal distribution of $Y_{i}$. We can thus consider a QQ plot of low through high quantiles of the $\textrm{Uniform}(0,1)$ distribution and the corresponding quantiles of the empirical CDF of $F(Y_{i},\widehat{\bm{\theta}}_{i})$s as a diagnostic tool. We refer to this tool as a Uniform-transformed QQ plot [hazra2020multivariate]. If the plot remains close to the $y=x$ line, we may conclude that $F$ provides a reasonable fit, in the sense that the model-implied distributions are consistent with the observed data. Conversely, systematic deviations highlight model misspecification, for example, a good fit in the bulk but a poor fit in the tails. If we compare two models, $F_{1}$ and $F_{2}$, we can calculate and compare the uniform-transformed QQ plots for both distributions.
Beyond visual inspection, one can quantify the departure from the $y=x$ line to yield a numerical goodness-of-fit measure. Two simple options are the root mean squared error and the mean absolute error, given by
$\text{RMSE}_{U}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(U_{(i)}-\frac{i}{n+1}\right)^{2}},~~~~\text{MAE}_{U}=\frac{1}{n}\sum_{i=1}^{n}\left|U_{(i)}-\frac{i}{n+1}\right|,$
respectively, where $U_{(i)}$ denotes the $i$th order statistic of $\{U_{i}\}$. This statistic measures the average squared distance between the empirical and theoretical quantiles, directly corresponding to deviations in the QQ plot. More formal alternatives include classical goodness-of-fit statistics such as the Kolmogorov–Smirnov [Kolmogorov1933, smirnov1948table], Cramér–von Mises [d1986goodness], or Anderson–Darling tests [anderson1954test], each emphasizing different aspects of deviation (maximum discrepancy, average discrepancy, or tail behavior). Thus, the uniform-transformed QQ plot provides a graphical diagnostic, and one can pair it with a scalar summary for model comparison. We will use this diagnostic tool later in Section 3 to validate the choice of gamma marginal distributions compared to alternatives such as the log-normal distribution.
@@ 2.5 Moran’s I for assessing spatial autocorrelation
Areal data often exhibit autocorrelation, meaning that observations at nearby, possibly neighboring, regions are more similar to each other than to those at distant ones. Moran’s I [moran1950notes] is a widely used measure for detecting spatial autocorrelation, and it is calculated as
@@
where $Y_{i}$ represents the observed value at region $i$, $w_{ij}$ is the spatial weight matrix defining neighborhood relationships in (3), $\bar{Y}$ is the mean value of $Y_{i}$, and $N$ is the total number of locations. Moran’s I ranges from 1-1 to +1+1, where positive values of Moran’s I suggest clustering, negative values indicate dispersion, and values near zero suggest spatial randomness. Under the null hypothesis of complete randomness, $\text{E}(I)=-1/(N-1)$ and $\text{Var}(I)$ can be calculated in terms of $w_{ij}$s, and we thus can obtain a $z$-score by centering and scaling. Finally, we obtain the $p$-value by assuming normality for the test statistic and decide on the spatial randomness accordingly.
@@ 2.6 Bayesian model comparison
To compare multiple Bayesian models with the same likelihood structure but different prior structures, we report Deviance information criteria [DIC, spiegelhalter2002bayesian] and Watanabe-Akaike information criteria [WAIC, watanabe2010asymptotic]. We calculate DIC as
$\textrm{DIC}=\bar{D}+p_{D}=D(\bm{Y}\mid\widehat{\bm{\theta}})+2p_{D},$
[…]
Further, we can compute WAIC, the widely applicable information criterion, as an alternative to DIC. Here, WAIC approximates $n$-fold (i.e., leave-one-out) cross-validation. Suppose we have $n$ independent observations (possibly vectors) $\bm{Y}_{i}$ for $i=1,\ldots,n$ and $\bm{\theta}$ is the parameter vector in the model (likelihood) $f(\cdot)$. Let $m_{i}$ and $v_{i}$ be the posterior mean and variance of $\log\left[f\left(\bm{Y}_{\bm{i}}\mid\bm{\theta}\right)\right]$. The effective model size is $p_{W}=\sum_{i=1}^{n}v_{i}$ and WAIC is given by
@@
The measure DIC is computationally simpler than WAIC but may be unreliable for hierarchical and complex models. At the same time, WAIC is more general and provides a better approximation of predictive performance, making it preferable in most Bayesian applications. [vehtari2017practical] recommend using WAIC or leave-one-out cross-validation over DIC for model comparison in modern Bayesian workflows.
@@ 3 Data description and exploratory analysis
Unlike the administrative state-wise boundaries, mainland India (excluding the Andaman and Nicobar Islands and the Lakshadweep Archipelago) comprises 34 meteorological subdivisions, which are determined based on meteorological homogeneity [guhathakurta2008trends]. We present the geographical boundaries of these 34 subdivisions in Figure 2. We obtain monsoon (July–September) total rainfall data (in mm) for these subdivisions, covering the years 1951–2014, from the Open Government Data (OGD) Platform, India (https://data.gov.in), along with shape files of the boundaries. Based on the rainfall data collected from 641 districts across India by the India Meteorological Department (IMD), [guhathakurta2008trends] computed monthly rainfall amounts for these districts by averaging the rainfall data from available stations within each district for each month. An area-weighted averaging method was then applied to derive subdivision-wise rainfall data from the district-wise data. [guhathakurta2011impact] provide a more detailed report on the dataset preparation. The rest of the paper focuses on analyzing this dataset, which we refer to as the “Indian rainfall dataset”.
Suppose we denote the monsoon total rainfall for the $i$th meteorological subdivision and $t$th year by $Y_{it}$, for $i=1,\ldots,n$ and $t=1,\ldots,T$ with $n=34$ and $T=64$. Given the extensive 6464–year observational period encompassing phases of global warming, significant trends across the years are plausible. The data are positive-valued; hence, we can model the marginal distributions using log-normal or gamma regression, for example, where we can model the mean as a function of the year. A log-normal distribution assumption is more prevalent in the general spatial or areal data analysis literature because it allows for a logarithmic transformation of the data, enabling model fitting using an ordinary or spatial linear regression framework, rather than a generalized ordinary or spatial linear regression framework. At the same time, as mentioned in Section 1, the gamma distribution assumption for the marginal distributions is the most popular in rainfall modeling literature.
@@
We first postulate that the data are independently distributed across years and marginally follow a gamma distribution. To model the space-time varying mean parameter, we write $\text{E}(Y_{it})=\mu_{it}$ with spatially varying shape parameter $a_{i}$. Hence, for every region index $i$, we assume
@@
where $a_{i}$ is the shape parameter and $\mu_{it}$ is the mean of the gamma distribution for region index $i$ and time point $t$, which is connected with a log-link function
@@
in which $\alpha_{i}$ and $\beta_{i}$ are spatially varying coefficients (SVCs). A competing log-normal model would be $Y^{*}_{it}=\log(Y_{it})\overset{\mathrm{ind}}{\sim}\text{Normal}(\mu_{it},\sigma^{2}_{i}),\ t=1,\dots,T$, where $\mu_{it}=\alpha_{i}^{*}+\beta_{i}^{*}t$ for SVCs $\alpha^{*}_{i}$, $\beta^{*}_{i}$, and variances $\sigma^{2}_{i}$. We separately fit the regression models for every region index $i$ using maximum likelihood estimates and compare their uniform-transformed QQ plots in the left panel of Figure 1. The lognormal model provides a better fit for low quantile levels, while the gamma model outperforms the lognormal model overall. Compared to the lognormal model, the gamma model provides approximately 7% lower $\textrm{RMSE}_{U}$ and $\textrm{MAE}_{U}$. For every subregion $i$, we present the uniform-transformed quantiles based on gamma regression in the right panel of Figure 1. Except for certain middle-range quantiles, the model fits the data reasonably well. In the rest of the paper, we subsequently consider the gamma model in (8) and (9).
@@
We further reparameterize the gamma model in (8) and (9) in terms of the standard shape-rate parameters $(a_{i},\lambda_{it})$, as this representation clarifies the role of each parameter. The time covariate is centered and scaled to $t^{*}$ to reduce correlation between the intercept and slope and improve estimation stability [gelman2008weakly].
$t^{*}=\frac{t-m_{t}}{s_{t}},\quad m_{t}=\frac{1}{T}\sum_{t=1}^{T}t,\quad s_{t}^{2}=\frac{1}{T}\sum_{t=1}^{T}t^{2}-m_{t}^{2}.$
With this transformation, we may rewrite the log-link in (9) as
$\log(\mu_{it})=\tilde{\alpha}_{i}+\tilde{\beta}_{i}t^{*},$
where $\tilde{\alpha}_{i}=\alpha_{i}+\beta_{i}m_{t}$, $\tilde{\beta}_{i}=\beta_{i}s_{t}$.
Under the standard shape-rate gamma parameterization, we have the mean $\mu_{it}={a_{i}}/{\lambda_{it}}$. Substituting the above expression of $\mu_{it}$ yields
$\lambda_{it}=\frac{a_{i}}{\mu_{it}}=a_{i}\exp[-\tilde{\alpha}_{i}]\exp[-\tilde{\beta}_{i}t^{*}].$
For interpretability, let $b_{i}=\exp[-\tilde{\alpha}_{i}]$ and $c_{i}=-\tilde{\beta}_{i}$. Then the rate can be written compactly as
$\lambda_{it}=a_{i}b_{i}\exp[c_{i}t^{*}].$ (10)
Here, $a_{i}$ controls the shape (or equivalently, the coefficient of variation), $b_{i}$ is an intercept-related parameter, and $c_{i}$ is a slope-related parameter that captures the climate change trend for subdivision $i$. With this reparameterization, the mean and variance of $Y_{it}$ follow as
$\begin{split}\text{E}(Y_{it})&=\frac{a_{i}}{\lambda_{it}}=b_{i}^{-1}\exp[-c_{i}t^{*}],\\ \text{Var}(Y_{it})&=\frac{a_{i}}{\lambda_{it}^{2}}=\frac{1}{a_{i}}(b_{i}^{-1}\exp[-c_{i}t^{*}])^{2}.\end{split}$ (11)
Thus, while $b_{i}$ shifts the overall rainfall level for region $i$, $c_{i}$ determines the temporal trend, and the coefficient of variation remains constant over time at $a_{i}^{-1/2}$ so that variance-to-mean ratios remain unchanged across regions. In this reparameterization, expressing the model in terms of $(a_{i},\lambda_{it})$ clarifies the relationship between the mean and variance, and simplifies expressions for the variance (as in (11)). This decomposition provides interpretable parameters that can be meaningfully compared across different climatic regions of India.
For exploratory analysis, we fit the model in (8) and (10) separately for every region $i$ using maximum likelihood estimation procedure and we present the estimates (MLEs) of $\bm{a}=(a_{1},\dots,a_{n})^{\intercal}$, $\bm{b}=(b_{1},\dots b_{n})^{\intercal}$ and $\bm{c}=(c_{1},\dots,c_{n})^{\intercal}$ in Figure 2. The standard errors of the estimates are low due to $T=64$ replications used for the analysis, and hence, any spatial pattern visible in Figure 2 can be conceptualized to build a flexible model for them. First, we notice that the MLEs of $a_{i}$s, $b_{i}$s, and $c_{i}$s are non-constant across space, motivating us to consider SVCs. Second, the estimates for nearby regions exhibit similarity, indicating the suitability of a CAR or an ICAR prior for the parameters.